# Boosting Latent Diffusion with Flow Matching

### v1

@@ 5 Conclusion
Our work introduces a novel and effective approach to high-resolution image synthesis, combining the generation diversity of Diffusion Models, the efficiency of Flow Matching, and the effectiveness of convolutional decoders. Strategically integrating Flow Matching models between a standard latent Diffusion model and the convolutional decoder enables a significant reduction in the computational cost of the generation process by letting the expensive Diffusion model operate at a lower resolution and up-scaling its outputs using an efficient Flow Matching model. Our Flow Matching model efficiently enhances the resolution of the latent space without compromising quality. Our approach complements DMs with their advancements and is orthogonal to their recent enhancements such as sampling acceleration and distillation techniques e.g., LCM [39]. This allows for mutual benefits between different approaches and ensures the smooth integration of our method into existing frameworks.
Supplementary Material
Figure 11: Samples from the Coupling Flow Matching model on top of Stable Diffusion V1.5 [48] using a classifier-free guidance scale of $7.5$. Samples are generated in latent space $64^{2}$ and up-sampled with CFM from $64^{2}$ to $128^{2}$. The resulting image has a resolution of $1024\times 1024$. Best viewed when zoomed-in.
We initiate our analysis by ablations on the flow matching model, systematically exploring its individual components and assessing their impact on the overall performance. Subsequently, we delve into a comprehensive examination of the combined effects of the flow matching and diffusion models. This examination considers both image quality and the efficiency gains achieved through this integration.
6 Quantitative Experimental Results
6.1 Discussion on Flow Matching
Effects of ODE Solver Steps on Image Quality We use the torchdiffeq [11] library for ordinary differential equation (ODE) solvers implemented in PyTorch. Tab. 4 demonstrates the ability of our flow matching model to produce strong results based on FID with as few as 10 function evaluations with additional speedups when compared against diffusion models. This highlights the efficiency and advantage of Flow Matching models in contrast to score-based models, which typically require more function evaluations due to its inherent stochasticity. Fig. 12 shows the average time needed for up-sampling a latent code from $32\times 32$ to $128\times 128$ on a single NVIDIA A100 GPU with batch-size of $4$ and Coupling Flow Matching. At 10 steps, the FID score plateaus, emphasizing the minimal gains from additional steps and empirically demonstrating that modeling direct trajectories accelerates sampling speed.
@@ 6.1 Discussion on Flow Matching
NFE SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ Time $(s)$downarrow
@@
Table 4: We ablate the CFM-200 model and compare different number of function evaluations (NFE) during inference for the Euler method, as well as the adaptive step-size Dormand-Prince solver. We evaluate the results on a subset of 1k samples of FacesHQ and present the total inference time on a single NVIDIA A100 GPU.
Figure 12: Average computation time on a single NVIDIA A100 GPU with batch-size $4$ for different number of function evaluations with the Euler method and the Dormand-Prince adaptive step-size ODE solver on the FacesHQ dataset.
We additionally illustrate the outcomes of image up-sampling for various number of function evaluations (NFE) in Fig. 13. An intriguing observation emerges when working with an exceptionally low NFE, which leads to higher PSNR and SSIM but with a notably increased blurriness. These results highly resemble the regression baseline model. With increasing number of steps images get sharper, with marginal to no perceptual differences between $10$ and $50$ NFEs. This again emphasizes the low number of function evaluations required to obtain reasonable high-resolution results from low-resolution conditioning information.
GT NFE=1 NFE=2 NFE=4 NFE=10 NFE=50
Figure 13: Sample quality of different ODE solvers and different number of function evaluations (NFE). From left to right, 1st column represents the ground truth, high-resolution images. From 2nd column on, we show the results for NFE $={1,2,4,10,50}$ with the Euler fixed step-size ODE solver.
6.2 DMs for lower Resolutions
To evaluate the performance of our approach, we perform flow matching on also low-dimensional latents. The standard SD V1.5 [48] was trained to work on latents with a dimensionality of $64^{2}$ pixels. After decoding the latent, this results in generated images with a resolution of $512^{2}$ pixels. Sampling latents of a different dimensionality is also possible but results in degraded performance, which is particularly pronounced for sampling lower-resolution latents. This deterioration can be partially mitigated by changing the scale of the self-attention layers from the standard $\sqrt{1/d}$ to $\sqrt{\log_{T}N/d}$ where $d$ is the inner attention dimensionality and $T$ and $N$ the number tokens during the training and inference phase respectively [22]. We use this to rescale the attention layers of SD V1.5 for $32^{2}$ latents. Additionally, we finetune it on images of LAION-Aesthetics V2 6+ [1] rescaled to $256^{2}$ pixels for one epoch, a learning rate of $1e{-5}$, and batch size of $256$.
We provide fine-tuned LDM metrics on the LAION dataset in Tab. 5. We found that FID pleataus with a classifier-free guidance scale at $7.0$. Consequently, we adopt this value for small resolution ($32^{2}$ in the latent space), text-guided image synthesis. We illustrate the uncurated samples for the upsampled images in Fig. 14.
CFG FID $\downarrow$ CLIP $\uparrow$
1 41.95 0.157
3 17.13 0.214
5 14.03 0.231
7 13.47 0.239
9 13.50 0.243
Table 5: Results of our fine-tuned Latent Diffusion Model on the LAION dataset for different Classifier-free guidance scales. [17]
Figure 14: Uncurated samples from the Coupling Flow Matching model on top of finetuned Stable Diffusion for lower resolution using a classifier-free guidance scale of $7.0$. Samples are generated in latent space $32^{2}$ and up-sampled with CFM from $32^{2}$ to $128^{2}$. The resulting images have a resolution of $1024\times 1024$.
Figure 15: Absolute inference time per model part (LDM or CFM) and FID for synthesising images with resolution $1024\times 1024$ for different number of function evaluations (NFE). Left axis refers to average inference time in seconds, right axis to $\log$ FID for $1$k samples. Dopri5 corresponds to the Dormand-Prince adaptive step-size solver. For LDM $128^{2}$ we do not use any CFM model.
7 Qualitative Experimental Results
Fig. 18 shows additional, zoomed in results of our FacesHQ Coupling Flow Matching model, where we up-sample the latent code of $256^{2}$ images from $32^{2}$ to $128^{2}$ with the Dormand-Prince ODE solver. In Fig. 19 and Fig. 20 we visualize more, non-curated results from our Coupling Flow Matching model on the FacesHQ dataset and the LHQ dataset respectively.
Fig. 17 show generated text-to-image samples in $2048^{2}$ px using the LDM together with CFM up-sampling in latent space with an intermediate resolution of $512^{2}$, produced by SD V1.5.
8 Architecture and Implementation Details
In our implementation of FM, we follow Tong et al. [60] and set the minimal variance either as $\sigma_{\min}=1e{-8}$ or $\sigma_{\min}=1e{-4}$. We employ the cosine noising schedule [43] for noise augmentation [19], and $t\in[1,999]$ resembles the corresponding timestep. We adopt the widely used U-Net architecture as implemented in the OpenAI GitHub repository [2]. Implementation details and parameterization for the UNet used for our Coupling Flow Matching models can be found in Tab. 8.
KL Autoencoder
We employ the KL-regularized autoencoder from Stable Diffusion V1.5 [48] to compress images to their respective latent representations. We also shed light on how well the pre-trained autoencoder from Stable Diffusion perform on smaller and larger resolution. We focus here on the autoencoder architecture that we employ for latent compression, since it is crucial to ensure that the latent space captures perceptual information at various resolution. Thanks to the convolutional nature, the autoencoder exhibits the capability to encode images at multiple scales. To systematically evaluate its effectiveness, we present results using metrics such as FID, PSNR, and SSIM for images of different sizes. The detailed metrics are presented in Tab. 6, and the FID is calculated with a representative sample size of $10,000$ images. Fig. 16 additionally shows visual results. Our analysis underscores the autoencoder’s proficiency across multiple resolutions. Despite potential shortcomings in low-resolution compression, we contend that the FM model can adeptly learn to transition from this distribution to the latent space of the high-resolution image, thereby mitigating deficiencies in low-resolution performance.
@@ 8 Architecture and Implementation Details
Figure 16: Comparison between reconstructed images (middle $256^{2}$ and right column $2048^{2}$) and their high-resolution original input (left column) using the pre-trained autoencoder. We can observe that the (i) pre-trained autoencoder is able to encode and decode images at different scales. (ii) It cannot reconstruct faces correctly in low resolution. (iii) The artifacts diminish with a higher resolution. Best viewed when zoomed in.
LAION-10k
Image Size SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
256 0.82 26.42 2.47
512 0.86 28.65 1.28
1024 0.88 30.72 0.84
2048 0.88 32.20 0.60
Table 6: Evaluation of the pre-trained autoencoder performance under different image sizes.
Big (306M) Small (113M)
Model SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
FM 0.59 22.90 11.22 0.59 22.84 10.90
CFM-400 0.59 22.72 6.30 0.59 22.87 6.38
CFM-200 0.60 22.98 4.92 0.60 23.02 5.20
Table 7: LHQ ablation for $256^{2}\rightarrow 1024^{2}$ super-resolution results.
FacesHQ Unsplash LHQ
Autoencoder $f$ 8 8 8
$z$-shape $4\times 64\times 64$ $4\times 64\times 64$ $4\times 64\times 64$
@@
Optimizer Adam Adam Adam
Batch size 96 768 128
Learning rate 1e-4 1e-4 1e-4
Table 8: Hyperparameters and number of parameters for our FMs and DMs. We use the exact same architecture for our FM and the SR3 [51] upsampling model to ensure a fair comparison. We trained the SR3 model with $1000$ diffusion steps and a cosine noise schedule [43].
Figure 17: Samples from the Coupling Flow Matching model on top of LCM-LoRA in 2k resolution. Samples are generated in latent space $64^{2}$ and up-sampled with CFM from $64^{2}$ to $256^{2}$. The resulting image has a resolution of $2048\times 2048$. Best viewed when zoomed-in.
GT REG LR FM
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 18: Qualitative comparison of the Flow Matching (FM) model to the upsampled low resolution image (LR) the regression model (REG) and the ground truth (GT) using zoomed-in patches. The FM patches are significantly sharper and display more fine-grained details.
Figure 19: Uncurated samples from the Coupling Flow Matching model on FacesHQ. Up-sampling in latent space from $32\times 32$ to $128\times 128$. Resulting image with resolution of $1024\times 1024$ (bottom) compared to the low-resolution image, up-sampled to $1024\times 1024$ (top). Best viewed when zoomed-in.
Figure 20: LHQ super-resolution samples. Left: low-resolution ground truth image bi-linearly up-sampled. Right: synthesized high resolution image. Best viewed when zoomed in.

### v2

@@ 5 Conclusion
Our work introduces a novel and effective approach to high-resolution image synthesis, combining the generation diversity of Diffusion Models, the efficiency of Flow Matching, and the effectiveness of convolutional decoders. Integrating Flow Matching models between a standard latent Diffusion model and the convolutional decoder enables a significant reduction in the computational cost of the generation process by letting the expensive Diffusion model operate at a lower resolution and up-scaling its outputs using an efficient Flow Matching model. Our Flow Matching model efficiently enhances the resolution of the latent space without compromising quality. Our approach complements DMs with their advancements and is orthogonal to their recent enhancements such as sampling acceleration and distillation techniques e.g., LCM [43]. This allows for mutual benefits between different approaches and ensures the smooth integration of our method into existing frameworks.
A Appendix for Boosting Latent Diffusion with Flow Matching
Figure 11: FID over training iterations for bilinear, nearest, and pixel-space upsampling of the low-resolution latent code on FacesHQ.
A.1 Degraded Image Super-Resolution
Originally, our model was designed to boost the resolution of existing diffusion models at a reduced cost. This differs from traditional super-resolution (SR) methods in two ways. First, they usually perform SR at lower resolutions, and second, they apply image degradation methods, e.g., compression artifacts, to obtain low-resolution images. Our model is not explicitly trained to be invariant to these degradations, but we can generalize to them. To support this claim, we additionally fine-tune our model on the Unsplash dataset, including image degradations following [69] on a 4×4$\times$4 × upsampling task from $128^{2}$ to $512^{2}$ pixels. We then perform zero-shot inference on the two common SR benchmark datasets DIV2K [3] and RealSR [10]. Tab. 5 quantitatively shows the comparison to other state-of-the-art SR methods, where our method achieves comparable or even superior results in FID and CLIP-IQA. Fig. 12 further compares the FID of our method against StableSR [66] in the low number of function evaluations (NFEs) regime. Our model excels particularly at low NFEs due to its straighter trajectories compared to the diffusion-based counterpart. Fig. 13 shows additional qualitative super-resolution results for mapping degraded low-resolution images to high-resolution images with our proposed CFM method.
Figure 12: In the low NFE regime we outperform StableSR [66] on FID at the $128^{2}\rightarrow 512^{2}$ px super-resolution task.
DIV2K [3] RealSR [10]
[…]
BSRGAN 0.63 24.60 44.22 0.52 0.77 23.69 141.28 0.50
RealERSGAN 0.64 24.33 37.64 0.52 0.76 25.69 135.18 0.44
DASR 0.63 24.50 49.16 0.50 0.77 27.02 132.63 0.31
LDM 0.58 23.36 36.21 0.62 0.72 25.49 132.32 0.59
StableSR 0.57 23.31 24.67 0.67 0.70 24.69 127.20 0.62
ResShift 0.62 24.69 36.01 0.61 0.74 26.31 142.81 0.55
CFM (ours) 0.52 21.63 29.76 0.73 0.65 23.31 125.88 0.67
Table 5: Quantitative comparison to other State-of-the-Art Super-Resolution models on two benchmark datasets. Comparison models are BSRGAN [74], RealERSGAN [68], DASR [34], LDM [53], StableSR [66], and ResShift [73].
A.2 Additional Visualizations
A.2.1 Image Synthesis at 1024 and 2048 Resolutions
We present additional samples generated at resolutions of 1k and 2k from our pipeline in Figs. 6, 15, 16 and 17. Our method produces high-resolution images with remarkable fidelity while maintaining fast inference times.
A.2.2 Super-Resolution
Fig. 13 shows qualitative results of our method applied to the image super-resolution task. Low-resolution images include degradations following [69] and we increase the resolution from $128^{2}$ to $512^{2}$ px. These results correspond to Tab. 5 from the main paper and showcase the generalizability of our model for super-resolution tasks.
Ground Truth Degraded Low-Res CFM Prediction
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Qualitative results for image super-resolution on degraded images from $128^{2}$ to $512^{2}$ px with our CFM model.
Figure 14: Samples decoded along the ODE trajectory for FacesHQ, LHQ, and LAION at $t\in\{0,0.25,0.5,0.75,1\}$, with a total number of function evaluations of 100. Best viewed when zoomed in.
Figure 15: Samples synthesized in $1024^{2}$ px. The comparison is highlighted on the top-left corners between the samples generated solely from DM and those generated from the combination between DM and CFM.
Figure 16: Chaining our models enables elevating the image resolution from $128^{2}$ to $2048^{2}$ px. The contrast before and after upsampling is presented in the right column, with the original low-resolution image positioned in the top-right corner for reference.
Figure 17: Upsampling results from $512^{2}$ to $2048^{2}$ px. The contrast before and after upsampling is presented in the right column, with the original low-resolution image positioned in the top-right corner for reference.
A.3 Additional Ablations
A.3.1 Up-Sampling Method
Our method requires matching resolutions between the source and target latent codes. We ablate three different options. The first two options involve directly upsampling the latent codes in the latent space, either via bilinear or nearest neighbor upsampling. The third method pixel space upsampling (PSU) first decodes the image into pixel space, then bilinearly upsamples the image, and encodes it back to latent space. This introduces some additional cost through the decoding and encoding operation, however, we found it to be negligible small.
Tab. 6 shows that pixel space upsampling (PSU) performs particularly well. Besides performing better overall, PSU also mitigates the problem of artifacts for low-resolution images. Fig. 18 and Fig. 19 show that the autoencoder introduces artifacts when encoding and decoding low-resolution images, e.g. of resolution $128^{2}$ and $256^{2}$ px. Increasing the resolution in pixel space and then encoding and decoding the image, avoids this issue. Since our primary goal is to speed up inference for image synthesis, using pixel space upsampling maintains most of the information present in the low-resolution latent code, which in turn ensures a well defined starting point for our Coupling Flow Matching module. This also makes the overall training more efficient as shown in
Original $\mathcal{D}(\mathcal{E}(x_{128}))$ $\mathcal{D}(\mathcal{E}(x_{128})\rightarrow 1024)$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 18: Left: Original image with $128^{2}$ px resolution, bilinearly upsampled to $1024^{2}$ px. Middle: $128^{2}$ px image encoded and decoded with the autoencoder, clearly showing artifacts. Right: $128^{2}$ px image bilinearly upsampled to $1024^{2}$ px in pixel space and then encoded and decoded with the autoencoder. Besides being blurry, the decoded image shows no artifacts.
FacesHQ LHQ
[…]
Bilinear 25.68 0.71 0.012 3.30 3.96 23.07 0.60 0.036 3.81 4.13
Nearest 25.55 0.71 0.013 3.32 3.77 22.81 0.59 0.038 4.03 4.67
Pixel Space 30.40 0.82 0.004 1.35 1.61 25.49 0.68 0.022 2.32 2.70
Table 6: Ablation of upsampling methods for our Coupling Flow Matching model in the latent space. The pixel space upsampling (PSU) method exhibits constantly better results than the latent space upsampling methods.
LAION-10k
Image Size SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
256 0.82 26.42 2.47
512 0.86 28.65 1.28
1024 0.88 30.72 0.84
2048 0.88 32.20 0.60
Table 7: Evaluation of the pre-trained autoencoder for different image resolutions.
A.3.2 Number of Function Evaluations
Tab. 8 shows the performance of our CFM model for different number of function evaluations. We evaluate the results on a subset of 1k samples of FacesHQ and present the total inference time on a single NVIDIA A100 GPU. We can clearly observe that our model achieves good performance already with as few as $10$ euler steps, indicating straight ODE trajectories.
@@ A.3.2 Number of Function Evaluations
NFE SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ Time (ms/img) $\downarrow$
@@
Table 8: We ablate the CFM-400 model and compare different numbers of function evaluations (NFE) during inference for the Euler method, as well as the adaptive step-size Dormand-Prince solver.
A.3.3 Autoencoder Resolutions
Even though the autoencoder from LDM [53] was trained on a fixed resolution, we find it to generalize well to images at different scales. This is shown quantitatively in Tab. 7, as well as qualitatively in Fig. 19.
Original $256^{2}$ $2048^{2}$
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ A.3.3 Autoencoder Resolutions
Figure 19: Comparison between reconstructed images and their high-resolution original input using the same pre-trained autoencoder. We can observe that the (i) pre-trained autoencoder can encode and decode images at different scales. (ii) It cannot reconstruct faces correctly in low resolution. (iii) The artifacts diminish with a higher resolution. Best viewed when zoomed in.
A.4 Training Details
The model architecture details for different datasets are provided in Tab. 9, and we employ a learning rate of $5\times 10^{-5}$. We precompute the image latents to enhance computational efficiency. For fairness, we train the diffusion model counterpart for upsampling using the same architecture, utilizing 1000 diffusion steps and a cosine diffusion schedule.
FacesHQ LHQ Unsplash
@@ A.4 Training Details
Batch size 96 128 768
Table 9: Hyperparameters and number of parameters for our Coupling Flow Matching module.
A.4.1 LDM $32^{2}$
In the main paper we use a Latent Diffusion Model that operates on a lower-dimensional latent space of $32^{2}\times 4$. The standard SD1.5 [53] was trained to work on latents with a dimensionality of $64^{2}$ pixels. After decoding the latent, this results in generated images with a resolution of $512^{2}$ pixels. Sampling latents of a different dimensionality is also possible but results in degraded performance, which is particularly pronounced for sampling lower-resolution latents. This deterioration can be partially mitigated by changing the scale of the self-attention layers from the standard $\sqrt{1/d}$ to $\sqrt{\log_{T}N/d}$ where $d$ is the inner attention dimensionality and $T$ and $N$ the number tokens during the training and inference phase respectively [23]. We use this to rescale the attention layers of SD-1.5 for $32^{2}$ latents. Additionally, we finetune it on images of LAION-Aesthetics V2 6+ [1] rescaled to $256^{2}$ pixels for one epoch, a learning rate of $1e{-5}$, and batch size of $256$.
In Tab. 10 we provide metrics for our fine-tuned Latent Diffusion Model on the LAION dataset. We find that FID plateaus with a classifier-free guidance scale at $7.0$. Consequently, we adopt this value for small resolution ($32^{2}$ in the latent space), text-guided image synthesis.
CFG scale FID $\downarrow$ CLIP $\uparrow$
1 41.95 0.157
3 17.13 0.214
5 14.03 0.231
7 13.47 0.239
9 13.50 0.243
Table 10: Results of our fine-tuned Latent Diffusion Model on the LAION dataset for different Classifier-free guidance (CFG) scales [18].